# Discriminative Entropy Clustering and its Relation to K-means and SVM

### v3

@@ 1 Introduction
$k_{\mu}(X):=\operatorname*{arg\,min}_{k}\,\|X-\mu_{k}\|$ $\sigma_{\bf v}(X):=\sigma({\bf v}^{\top}X)\equiv\text{soft-max}({\bf v}^{\top}X)$
@@
Figure 1: Entropy clustering vs. K-means - binary example ($K=2$) for 2D data $\{X_{i}\}$ comparing two linear methods of similar parametric complexity: (a) $K$-means $\mu_{k}\in{\cal R}^{2}$ and (b) entropy clustering (5) with a linear model (3) defined by $K$-column matrix ${\bf v}=[{\bf v}_{k}]$ with linear discriminants ${\bf v}_{k}\in{\cal R}^{2+1}$ (incl. bias). Red and green colors in (a) and (b) illustrate decision/prediction functions, $k_{\mu}(X):=\operatorname*{arg\,min}_{k}\,\|X-\mu_{k}\|$ and $\sigma_{\bf v}(X):=\sigma({\bf v}^{\top}X)$, corresponding to the optimal parameters $\mu$ and ${\bf v}$ minimizing two losses: (a) compactness or variance of clusters $\sum_{i}\|X_{i}-\mu_{k_{i}}\|^{2}$ where $k_{i}=k_{\mu}(X_{i})$, and (b) decisiveness and fairness $\overline{H(\sigma)}-H(\bar{\sigma})$, see (5), where $\sigma_{i}=\sigma({\bf v}^{\top}X_{i})$. Color transparency in (b) visualizes “soft” decisions $\sigma({\bf v}^{\top}X)$; the linear boundary “blur” is proportional to $\frac{1}{\|v\|}$. Unlike low-variance (a), the optimal clusters in (b) have the maximum margin among all fair/balanced solutions, assuming “infinitesimal” norm regularization $\|{\bf v}\|^{2}$ discussed in Sec.2.2.
@@ 2.2.1 Overview of max-margin clustering
As typical in SVM, our formal theories on the max-margin property of the regularized decisiveness (13) in Sections 2.2.3-2.2.4 are mainly focused on the binary linear models, though multi-class extensions are possible, see Sec. 2.2.5. Below we informally preview our max-margin property claims about (13) providing some intuition that may be helpful due to differences with standard margin maximization in SVM.
For example, basic SVM assumes linearly separable data in a fully supervised setting. Soft SVM extension allows a trade-off between margin maximization and separation violations. In contrast, linear separability is irrelevant in clustering as there are no ground-truth labels to be violated. Yet, the “gap” can be measured between any pair of clusters, see Fig. 3(a) and (b). Thus, the largest margin solution among all feasible (fair) linear clusterings is well-defined, see Fig. 3(a).
@@ 2.2.5 Generalization for $K>2$
Figure 8: Multi-label clustering: K-means vs. entropy clustering. Consistently with Theorem 4, loss (14) produces max-margin clustering, as in Definitions 3-4, satisfying the fairness.
@@ 2.2.6 Discussion: entropy clustering vs SVM clustering
where, unlike supervised SVM [42], binary targets $t\in\{\pm 1\}$ are optimization variables here. They are estimated jointly with the linear model parameters ${\bf v}$. Following our terminology from Section 1.4, objective (2.2.6) is a self-labeling loss with binary pseudo-labels $t$. It resembles (8) where hard constraints represent fairness and pseudo-labels are discrete/hard. One can also relate the hinge loss in (2.2.6) to a truncated variant of decisiveness [^5] Minimizing out $t$ in the hinge loss (2.2.6) gives $\max\{0,1-|{\bf v}^{\top}X|\}$. encouraging linear discriminant ${\bf v}$ such that $|{\bf v}^{\top}X|\geq 1$ for all data points $X$.
@@ 3 Our self-labeling algorithm
The theories in the previous section improve the general understanding of the properties of discriminative entropy clustering. In particular, they show certain conditions when the linear classifier norm regularization leads to margin maximization. Note that the generic weight decay implicitly represents the norm regularization. Still, it may not satisfy these conditions as it often includes the bias parameter, and its strength may not be set correctly, e.g. to a tiny $\gamma$ value.
This section addresses other limitations of prior entropy clustering algorithms. We focus on self-labeling (Sec.1.4) and show that the standard cross-entropy formulation of decisiveness is sensitive to pseudo-label errors. Section 3.1 argues that the reverse cross-entropy is more robust to label noise. We also propose strong fairness. Section 3.2 derives an efficient EM algorithm for minimizing our loss w.r.t. pseudo-labels, which is a critical step of our self-labeling algorithm.
@@ 3.1 Self-labeling surrogate loss formulation
which has one notable difference from the standard self-labeling surrogate losses (8) and (9) from [6, 7]: the cross-entropy term has the reverse order of the arguments $y$ and $\sigma$. Note that the standard order is motivated in the supervised settings where $y$ is a true one-hot target and $\sigma$ is the estimated distribution. However, this motivation does not work when both $\sigma$ and $y$ are estimated soft distributions. Moreover, Figure 10(b) argues that the reverse cross-entropy can improve the robustness of predictions $\sigma$ to errors in estimated pseudo-labels $y$, as confirmed by our network training experiment in Figure 11. The reversal also works for estimating soft pseudo-labels $y$ as the second position in cross-entropy is standard for an “estimated” distribution.
@@
Due to the opposite order in both cross-entropy and KL terms compared to (9), this loss represents stronger constraints for $y$ and $\bar{y}$. The next Section discusses an efficient EM algorithm for estimating pseudo-labels based on (42).
Figure 11: Robustness to noisy labels: reverse $H(\sigma,y)$ vs standard cross-entropy $H(y,\sigma)$. We train ResNet-18 on fully-supervised Natural Scene dataset [44] where we corrupted some labels. The horizontal axis shows the corruption level, i.e. percentage $\eta$ of training images where correct ground truth labels were replaced by a random label. We use soft target distributions $\tilde{y}=\eta*u+(1-\eta)*y$ that is a mixture of one-hot distribution $y$ for the observed corrupt label and the uniform distribution $u$, as in [35]. The vertical axis shows the test accuracy. Reverse cross-entropy improves robustness to high labeling errors.
@@ 3.2 Our EM algorithm for pseudo-labels
Minimization of a self-supervised loss w.r.t pseudo-labels $y$ for given predictions $\sigma$ is a critical operation in iterative self-labeling techniques [6, 7], see Sec.1.4. Besides well-motivated numerical properties of our new loss (42), in practice it also matters that it has an efficient solver for pseudo-labels. While (42) is convex w.r.t. $y$, optimization is done over a probability simplex and a good practical solver is not a given. Note that $H(\sigma,y)$ works as a log barrier for the constraint $y\in\Delta^{K}$. This could be problematic for the first-order methods, but a basic Newton’s method is a good match, e.g. [45]. The overall convergence rate of such second-order methods is fast, but computing the Hessian’s inverse is costly. Instead, we derive a more efficient expectation-maximization (EM) algorithm, see Table I, as discussed below.
@@ 3.3 Discussion
Here we briefly discuss the differences between our method for (42) and the algorithm for (12) in [7]. Their loss is convex with respect to ${\bf v}$. In contrast, the reverse cross-entropy in our loss (42) is non-convex as a natural consequence of our focus on robustness to pseudo-label errors. However, both losses are non-convex w.r.t. parameters ${\bf w}$ of the interior network layers responsible for representation $f_{\bf w}(X)$ in (4). As for the optimization w.r.t. the pseudo-labels $y$, both losses are convex. Our EM algorithm converges to the global optima for $y$, while they use an approximate closed-form solution for $y$, but it does not have optimality guarantees. Both self-labeling algorithms iteratively optimize the corresponding surrogate losses over pseudo-labels $y$ and model parameters ${\bf v}$, ${\bf w}$.
@@ 4 Experimental results
We test our approach on standard benchmark datasets with different network architectures. We also provide a comparison of different losses under semi-supervised settings. All the results are reproduced using either public codes or our own implementation under the same experimental settings for fair comparison.
Dataset For the clustering problem, we use four standard benchmark datasets: MNIST [47], CIFAR10/100 [48] and STL10 [49]. We follow [5] to use the whole dataset for training and testing unless otherwise specified. As for the semi-supervised setting, we conduct experiments on CIFAR10 and STL10. We split the data into training and test sets as suggested by the official instructions of the datasets.
Evaluation As for the evaluation of clustering, we set the number of clusters to the number of ground-truth category labels. To calculate the accuracy, we use the standard Hungarian algorithm [50] to find the best one-to-one mapping between predicted clusters and ground-truth labels. We don’t need this matching step if we use other metrics, i.e. NMI, and ARI. For the evaluation of semi-supervised classification, we directly calculate the accuracy of the test set without the mapping step.
Implementation details For clustering with fixed features, we set $\lambda$ in our loss to 100. The learning rate of the stochastic gradient descent is set to 0.1 and training takes 10 epochs. The batch size was set to 250. The coefficients for the $l_{2}$ norm of the weights are set to 0.001, 0.02, 0.009, and 0.02 for MNIST, CIFAR10, CIFAR100, and STL10 respectively. We also tuned the hyperparameters for all other methods.
@@
4.1 Effect of $\alpha$ in Renyi entropies for clustering
In Section 2.2, we provide theoretical results that any $\alpha$ in Renyi entropies leads to the maximum margin under certain conditions. However, it is unclear how different $\alpha$ affects the results in real-world settings. Here, we use the standard STL10 dataset to test the clustering performance for different $\alpha$ following the experimental settings from [52].
@@ 4.1 Effect of $\alpha$ in Renyi entropies for clustering
TABLE II: Effects of different $\alpha$ in Renyi entropy as the decisiveness term.
We tuned the hyperparameters for each column and used the stochastic gradient descent as the optimizer. While our self-labeling approach can give a performance boost for $\alpha=1$, it does not work well empirically for other $\alpha$. As one can easily check, the subproblems in the self-labeling formulation for $\alpha\neq 1$ can not be easily optimized. From the results in Table II, we can see that larger $\alpha$ gives better results. Note that larger $\alpha$ induces smaller gradients as the network prediction gets more confident. If the network prediction is wrong, there is a chance for correction due to small gradients. On the contrary, the smaller $\alpha$ has larger gradients as network predictions become more confident, which creates strong local minima that are hard to get out of.
@@ 5 Conclusions
Our paper clarified several important conceptual properties of the general discriminative entropy clustering methodology. We formally proved that linear entropy clustering has a maximum margin property, establishing a conceptual relation between it and SVM clustering. Unlike prior work on discriminative entropy clustering, we show that classifier norm regularization is important for margin maximization. We also provided counterexamples, disproving a theoretical claim in a recent TPAMI paper stating the equivalence between variance clustering (K-means) and entropy clustering.
We also discussed several limitations of the existing self-labeling formulations of entropy clustering and proposed a new loss addressing such limitations. In particular, we replace the standard (forward) cross-entropy by the reverse cross-entropy that we show is significantly more robust to errors in estimated soft pseudo-labels. Our loss also uses a strong formulation of the fairness constraint motivated by a zero-avoiding version of KL divergence. Moreover, we designed an efficient EM algorithm minimizing our loss w.r.t. pseudo-labels; it is significantly faster than standard alternatives, e.g. Newton’s method. Our empirical results improved the state-of-the-art on many standard benchmarks for end-to-end deep clustering methods.
Appendix A Proof
Lemma 1.
Given fixed $\sigma_{i}\in\Delta^{K}$ where $i\in\{1,...,M\}$ and $u\in\Delta^{K}$, the objective
$E(y)=-\frac{\beta}{M}\sum_{i}\sum_{k}\sigma_{i}^{k}\ln{y_{i}^{k}}-\lambda\sum_{k}u_{k}\ln{\frac{\sum_{i}y_{i}^{k}}{M}}$
is convex for $y$, where $y_{i}\in\Delta^{K}$.
Proof.  First, we rewrite $E(y)$
$E(y)$ $=-\sum_{k}\left(\frac{\beta}{M}\sum_{i}\sigma_{i}^{k}\ln{y_{i}^{k}}+\lambda u_{k}\ln{\frac{\sum_{i}y_{i}^{k}}{M}}\right)$
$:=-\sum_{k}f_{k}(y^{k})$ (b)
Next, we prove that $f_{k}:R_{(0,1)}^{M}\to R$ is concave based on the definition of concavity[43] for any $k\in\{1,...,K\}$. Considering $x=(1-\alpha)x_{1}+\alpha x_{2}$ where $x_{1},x_{2}\in R_{(0,1)}^{M}$ and $\alpha\in[0,1]$, we have
[…]
$\geq\frac{\beta}{M}\sum_{i}(1-\alpha)\sigma_{i}^{k}\ln{x_{1i}}+\alpha\sigma_{i}^{k}\ln{x_{2i}}$
$\quad+\lambda u_{k}\left((1-\alpha)\ln{\frac{\sum_{i}x_{1i}}{M}}+\alpha\ln{\frac{\sum_{i}x_{2i}}{M}}\right)$
$=(1-\alpha)f_{k}(x_{1})+\alpha f_{k}(x_{2})$
The inequality uses Jensen’s inequality. Now that $f_{k}$ is proved to be concave, $-f_{k}$ will be convex. Then $E(y)$ can be easily proved to be convex using the definition of convexity with the similar steps above.

### v4

@@ 1 Introduction
$k_{\mu}(X):=\arg\min_{k}\,\|X-\mu_{k}\|$ $\sigma_{\bf v}(X):=\sigma({\bf v}^{\top}X)\equiv\text{soft-max}({\bf v}^{\top}X)$
@@
Figure 1: Entropy clustering vs. K-means - binary example ($K=2$) for 2D data $\{X_{i}\}$ comparing two linear methods of similar parametric complexity: (a) $K$-means $\mu_{k}\in{\cal R}^{2}$ and (b) entropy clustering (5) with a linear model (3) defined by $K$-column matrix ${\bf v}=[{\bf v}_{k}]$ with linear discriminants ${\bf v}_{k}\in{\cal R}^{2+1}$ (incl. bias). Red and green colors in (a) and (b) illustrate decision/prediction functions, $k_{\mu}(X):=\arg\min_{k}\,\|X-\mu_{k}\|$ and $\sigma_{\bf v}(X):=\sigma({\bf v}^{\top}X)$, corresponding to the optimal parameters $\mu$ and ${\bf v}$ minimizing two losses: (a) compactness or variance of clusters $\sum_{i}\|X_{i}-\mu_{k_{i}}\|^{2}$ where $k_{i}=k_{\mu}(X_{i})$, and (b) decisiveness and fairness $\overline{H(\sigma)}-H(\bar{\sigma})$, see (5), where $\sigma_{i}=\sigma({\bf v}^{\top}X_{i})$. Color transparency in (b) visualizes “soft” decisions $\sigma({\bf v}^{\top}X)$; the linear boundary “blur” is proportional to $\frac{1}{\|v\|}$. Unlike low-variance (a), the optimal clusters in (b) have the maximum margin among all fair/balanced solutions, assuming “infinitesimal” norm regularization $\|{\bf v}\|^{2}$ discussed in Sec.2.2.
@@ 2.2.1 Overview of max-margin clustering
As typical in SVM, our formal theories on the max-margin property of the regularized decisiveness (13) in Sections 2.2.3-2.2.4 are focused on the binary linear models, but some multi-class extensions are discussed in Sec. 2.2.5. Below we informally preview our max-margin property claims about (13) providing some intuition that may be helpful due to differences with standard margin maximization in SVM.
For example, basic SVM assumes linearly separable fully-labeled data. Soft SVM extension allows a trade-off between margin maximization and linear-separation violations. In clustering, however, any data can be linearly partitioned in many different ways as there are no ground-truth labels to be violated. The “gap” can be measured for any pair of such clusters, see Figs. 3(a,b). Thus, the largest margin solution among all feasible (e.g. fair) linear clusterings is well-defined.
@@ 2.2.5 Generalization for $K>2$
Figure 8: Multi-label clustering: K-means vs. entropy clustering. Consistently with Theorem 4, loss (14) produces max-margin clustering (b), as in Definitions 3-4, satisfying the fairness. The margin is tight only between the red and blue clusters.
@@ 2.2.6 Discussion: entropy clustering vs SVM clustering
where, unlike supervised SVM [42], binary targets $t\in\{\pm 1\}$ are optimization variables jointly estimated with the linear model parameters ${\bf v}$. Following our terminology from Section 1.4, objective (2.2.6) is a self-labeling loss with binary pseudo-labels $t$. It resembles (8) where hard constraints represent fairness and estimated pseudo-labels are discrete/hard. One can also relate the hinge loss in (2.2.6) to a truncated variant of decisiveness [^5] Minimizing out $t$ in the hinge loss (2.2.6) gives $\max\{0,1-|{\bf v}^{\top}X|\}$. encouraging linear discriminant ${\bf v}$ such that $|{\bf v}^{\top}X|\geq 1$ for all data points $X$.
@@ 3 Our self-labeling algorithm
The theories above improve the general understanding of discriminative entropy clustering and show certain conditions when the linear classifier norm regularization leads to margin maximization. Note that the generic weight decay implicitly represents the norm regularization $\gamma\|{\bf v}\|^{2}$, but it may not satisfy these conditions. For example, it often includes the bias parameter, and its strength $\gamma$ may not be set correctly.
This section addresses other limitations of prior entropy clustering algorithms in the context of Shannon’s entropy $H=R_{1}$ as in (14). We focus on self-labeling (Sec.1.4) and show that the standard cross-entropy formulation of decisiveness is sensitive to pseudo-label errors. Section 3.1 argues that the reverse cross-entropy is more robust to label noise. We also propose strong fairness. Section 3.2 derives an efficient EM algorithm for minimizing our loss w.r.t. pseudo-labels, which is a critical step of our self-labeling algorithm.
@@ 3.1 Self-labeling surrogate loss formulation
which has one notable difference from the standard self-labeling surrogate losses (8) and (9) from [6, 7]: the cross-entropy term reverses the order of its arguments $y$ and $\sigma$. The standard order is motivated in the supervised settings where $y$ is a true one-hot target and $\sigma$ is the estimated distribution. However, this motivation is questionable when both $\sigma$ and $y$ are estimated soft distributions. Moreover, Figure 10(b) argues that the reverse cross-entropy can improve the robustness of network predictions $\sigma$ to errors in estimated pseudo-labels $y$, as confirmed by our training experiments in Figure 11. The reversal also works for estimating soft pseudo-labels $y$ as the second argument in the cross-entropy function is standard for an “estimated” distribution.
@@
Due to the argument reversal in the cross-entropy and KL terms, our loss (42) represents stronger constraints for $y$ and $\bar{y}$ when compared to (9). However, besides well-motivated numerical properties, it also matters that (42) has an efficient solver for pseudo-labels discussed in the next Section.
Figure 11: Robustness to noisy labels: reverse $H(\sigma,y)$ vs standard cross-entropy $H(y,\sigma)$. We train ResNet-18 on fully-supervised Natural Scene dataset [44] where we corrupted some labels. The horizontal axis shows the corruption level, i.e. percentage $\eta$ of training images where ground truth labels were randomly replaced. We use true posteriors as soft target distributions $\tilde{y}=\eta*u+(1-\eta)*y$ that is a mixture of one-hot distribution $y$ for the observed corrupt label and the uniform distribution $u$, as in [35]. The test accuracy shows that the reverse cross-entropy is more robust to labeling errors.
@@ 3.2 Our EM algorithm for pseudo-labels
Efficient minimization of a self-supervised loss w.r.t pseudo-labels $y$ for given predictions $\sigma$ is a critical component of all iterative self-labeling techniques [6, 7], see Sec.1.4. While our loss (42) is convex w.r.t. $y$, optimization is done over the probability simplex and a good practical solver is not a given. Note that $H(\sigma,y)$ works as a log barrier for the constraint $y\in\Delta^{K}$. This could be problematic for the first-order methods, but a basic Newton’s method is a good match, e.g. [45]. The overall convergence rate of such second-order methods is fast, but computing the Hessian’s inverse is costly. Below we derive an expectation-maximization (EM) algorithm that is more efficient, see Table I.
@@ 3.3 Discussion
Here we briefly compare our optimization method for (42) and the algorithm for (12) in [7]. Their loss is convex with respect to ${\bf v}$. In contrast, the reverse cross-entropy in our loss (42) is non-convex as a natural consequence of our focus on robustness to pseudo-label errors. However, both losses are non-convex w.r.t. parameters ${\bf w}$ of the interior network layers responsible for representation $f_{\bf w}(X)$ in (4). As for the optimization w.r.t. the pseudo-labels $y$, both losses are convex. Our EM algorithm converges to the global optima for $y$, while they use an approximate closed-form solution for $y$, but it does not have optimality guarantees. Both self-labeling algorithms iteratively optimize the corresponding surrogate losses over pseudo-labels $y$ and model parameters ${\bf v}$, ${\bf w}$.
@@ 4 Experimental results
We test our approach to clustering on several standard datasets using different network architectures. We also evaluate different losses in semi-supervised settings. All the results for prior work are reproduced using either the corresponding public code or our implementation using the same experimental settings.
Datasets: For the clustering problem, we use four standard benchmark datasets: MNIST [47], CIFAR10/100 [48] and STL10 [49]. We follow [5] to use the whole dataset for training and testing unless otherwise specified. As for the semi-supervised setting, we conduct experiments on CIFAR10 and STL10. We split the data into training and test sets as suggested by the official instructions of the datasets.
Evaluation: To evaluate the quality of clustering, we set the number of clusters to the number of ground-truth category labels. To calculate the accuracy, we use the standard Hungarian algorithm [50] to find the best one-to-one mapping between predicted clusters and ground-truth labels. We don’t need this matching step if we use other metrics, i.e. NMI, and ARI. For the evaluation of semi-supervised classification, we directly calculate the accuracy of the test set without the mapping step.
Implementation details: For clustering with fixed features, we set $\lambda$ in our loss to 100. The learning rate of the stochastic gradient descent is set to 0.1 and training takes 10 epochs. The batch size was set to 250. The coefficients for the $l_{2}$ norm of the weights are set to 0.001, 0.02, 0.009, and 0.02 for MNIST, CIFAR10, CIFAR100, and STL10 respectively. We also tuned the hyperparameters for all other methods.
@@
4.1 Effect of Renyi entropy order $\alpha$
Our results in Section 2.2 state that the regularized decisiveness (15) using Renyi entropy of any order $\alpha>0$ leads to margin maximization under certain conditions, e.g. linear model and convergence of the regularization constant $\gamma$ to zero. However, such conditions may not be satisfied in practice. For example, only fixed values of $\gamma$ can be evaluated. Also, unsurprisingly, better results are typically obtained by non-linear network models where a linear classifier head is trained jointly with fine-tuning the representation layers. The corresponding non-convex losses may also be harder or easier to optimize depending on $\alpha$.
It follows that different values of $\alpha$ may affect the practical results. However, our self-labeling algorithm in Section 3 assumes $\alpha=1$ and uses the closed-form EM steps applicable only to the Shannon’s decisiveness. To evaluate Renyi decisiveness for various $\alpha>0$, we apply the standard stochastic gradient descent to a combination of (15) with Shannon’s fairness
$\gamma\|{\bf v}\|^{2}\;+\;\overline{R_{\alpha}(\sigma)}\;-\;H(\overline{\sigma}).$ (48)
We use the standard STL10 dataset following the experimental setting for clustering from [52]. All hyperparameters are separately tuned in each column.
@@ 4.1 Effect of Renyi entropy order $\alpha$
TABLE II: Effect of order $\alpha$ in Renyi decisiveness (15). The results use ResNet18 trained via standard stochastic gradient descent for loss (48). ACC for $\alpha=1$ is the same as for IMSAT [3] in Tab.V (STL10).
Table II suggests that larger values of $\alpha$ may lead to better empirical results. Intuitively, this could be explained as follows. Smaller $\alpha$ induces larger gradients for the Renyi decisiveness, see Figure 5, as the network prediction becomes more confident during training. This creates strong local minima for the network parameters that are hard to escape.
The results in Table II encourage the development of an efficient pseudo-labeling algorithm for $\alpha>1$, which is left for future work. The experiments in the following Sections use our efficient EM-based algorithm from Section 3 assuming Shannon’s entropy, i.e. $\alpha=1$. Note that our self-labeling optimization algorithm significantly outperforms the standard stochastic gradient descent for $\alpha=1$, see "our" vs IMSAT in Table V.
@@ 5 Conclusions
Our paper clarified several important conceptual properties of the general discriminative entropy clustering methodology. We formally proved that linear entropy clustering has a maximum margin property, establishing a conceptual relation to SVM clustering. Unlike prior work on discriminative entropy clustering, we show that classifier norm regularization is important for margin maximization. We also provided counterexamples disproving a recent theoretical claim of the equivalence between variance clustering (soft K-means) and linear entropy clustering (linear softmax classifier optimizing decisiveness and fairness). We juxtapose the two methods as generative and discriminative approaches to MI-based clustering [1].
We also discussed several limitations of the existing self-labeling formulations of discriminative entropy clustering and proposed a new loss addressing such limitations. In particular, we replace the standard (forward) cross-entropy by the reverse cross-entropy that we show is significantly more robust to errors in estimated soft pseudo-labels. Our loss also uses a strong formulation of the fairness constraint motivated by a zero-avoiding version of KL divergence. We designed an efficient EM algorithm for minimizing our loss w.r.t. pseudo-labels; it is significantly faster than standard alternatives, e.g. Newton’s method. Our empirical results improved the state-of-the-art on many standard benchmarks for end-to-end deep clustering methods.